kacchanforlife:

If the circumference of a circle is 30π in, find the area of the circle.

kacchanforlife:

@oliver69

OLIVER69:

Well you need to know the formula which is A = π r^2. But first you need to find the radius from the circumference, which is r = c / (2 * π) which if we plugged 30 it would be r = 30 / (2 * π ) Once you solve that you should get 4.77464829276, which is rounded to 4.77 Then you can plug that into the equation to find the area. A = π (4.77)^2 A = 71.4803434879 A = 71.5 inches